Grade 11Physical Chemistry

the value of kp for the equillibrium of the reaction N2O4 gives 2NO2 is 2 calculate % dissosciation of N2O4 at a pressure of 0.5 atm?

younus jalil
11 Years agoGrade 11

1 Answer

Sunil Kumar FP
11 Years ago
we have the equation as
N2O4 ---2NO2
Kp=(pNO2)^2/N2O4
let the %decomposition be x
initial pressure of N2O4 –.5
final pressure of N2O4 =.5-.5x
final pressure of NO2=.5x*2 =x
According to the problem
Kp=(x)^2/(.5-.5x)
2=x^2/.5(1-x)
4=2x^2/(1-x)
2-2x=x^2
x^2+2x-2 =0
x=.732 or 73.2%